OpenStudy (decarr432):

given f(x)= the square root of 9x and g(x)= 1 over x-1 Answer's given A. -5 B. 5 C. 1 D. -6

OpenStudy (decarr432):

I think I understand how to do it but I'm not to sure

OpenStudy (decarr432):

@mathmale

OpenStudy (decarr432):

@Jaynator495 @DanJS

OpenStudy (mathmale):

@Decarr432: I don't think you've included all of the directions for this problem. Please double check what you've typed in against the original problem statement. Thank you.

OpenStudy (decarr432):

Forgot one think lol Which is the domain of f o g

OpenStudy (decarr432):

So find f(x) first correct?

OpenStudy (mathmale):

f(x) is already given, so there's nothing to find. I'd prefer you write this function as\[f(x)=\sqrt{9x}\] and g(x) as\[g(x)=\frac{ 1 }{ x-1 }\]

OpenStudy (decarr432):

Okay, how do I find domain?

OpenStudy (astrophysics):

You want f(g(x)) as requested in your question

OpenStudy (astrophysics):

Can you set that up

OpenStudy (mathmale):

You are to answer the question: Which is the domain of f o g? You should form the actual composition f o g now (although in the future you may be able to do it upon inspection).

OpenStudy (mathmale):

f(x) and g(x) are given. Can you now form the composition (f o g)(x)? Follow exactlyl the same pattern you and I used before.

OpenStudy (decarr432):

Yes, and okay but how do I set it up with the g(x) being 1/x-1

OpenStudy (mathmale):

\[g(x)=\frac{ 1 }{ x-1 }\]

OpenStudy (decarr432):

I cant use the same tools as when answering a question I didn't ask buddy

OpenStudy (mathmale):

I use Equation Editor for that. If you don't know how to use Eq. Editor, you could always DRAW your equations and expressions, using the Draw utility (below).